OpenStudy (saad560):

What is the value of a? 5/6(a−8)=7

OpenStudy (saad560):

@563blackghost

OpenStudy (jcreamer):

Start by distributing the 6, so 5\[5\div6a -48=7\]

OpenStudy (saad560):

wait 5/6 is a fraction

563blackghost (563blackghost):

Wait is it... \(\huge{\frac{5}{6}(a-8)=7}\) ?

OpenStudy (saad560):

yes

563blackghost (563blackghost):

Well that clears the confusion :)

563blackghost (563blackghost):

First distribute `5\6` into the paranthesis.... \(\huge{((\frac{5}{6} \times a - (8 \times \frac{5}{6}))=7}\) What would that equal?

OpenStudy (jcreamer):

Ok the fraction would simplify as \[\frac{ 5(a-8) }{ 6 }=7\]

OpenStudy (saad560):

5a/6-20/3

OpenStudy (benji):

You could also get rid of the fraction by multiplying both sides by the reciprocal, I like to do that because I hate fractions...

563blackghost (563blackghost):

Correct :) \(\huge{\frac{5}{6}a-\frac{20}{3}=7}\) Now we would need to get like terms to one side....so we would add `20/3` on both sides... \(\huge{\frac{5}{6}a(-\frac{20}{3}+\frac{20}{3})=7+\frac{20}{3}}\) What would that equal?

OpenStudy (saad560):

5/6a=7 20/3

OpenStudy (saad560):

@563blackghost

563blackghost (563blackghost):

Well lets simplify that fraction....We multiply with the denominator by 7 and then add 20 to it to get the fraction....so it would be... \(\huge{\frac{5}{6}a=\frac{41}{3}}\)

OpenStudy (saad560):

ok

563blackghost (563blackghost):

Now we need to isolate `a` so we would divide by `5/6`.... \(\huge{(\frac{5}{6} \div \frac{5}{6})a=\frac{41}{3} \div \frac{5}{6}}\) What would a equal?

OpenStudy (saad560):

i need answer have to leave soon

OpenStudy (saad560):

@563blackghost i actualy have to go

OpenStudy (saad560):

actually*

563blackghost (563blackghost):

That equation will give you the answer I cant give it away sorry :(

OpenStudy (saad560):

16.4 i actually have to go i am serious @563blackghost

563blackghost (563blackghost):

I as well that is correct but its best to have it in fraction form so the answer would be \(\LARGE{\frac{82}{5}}\)